A random variable follows a binomial distribution with a probability of success equal to 0.64. For a sample size of n=9​, find the values below.

a. the probability of exactly 3 successes

b. the probability of 6 or more successes

c. the probability of exactly 9 successes

d. the expected value of the random variable
Answer #1

Solution p=0.64 n = 9 9-3 a) p (@=3) = 96₂ (0.64² (1-0.64) 10.0479 4.79% 9-6 b) p (a>,6)=f(x=6) + p(x=7) + p(x-8) + P (x=9) 9